OpenStudy (anonymous):

If gcd(x,y) = d, express gcd (18x + 3y, 3x) in terms of d. Justify your answer.

OpenStudy (anonymous):

For every a and b you have gcd(a, b) = gcd(a, b - a), you can apply this to your problem and you get gcd(18x + 3y, 3x) = gcd(3y, 3x) by substracting repeatedly 3x from the left side, and gcd(3y, 3x) = 3*gcd(y, x) = 3d

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

No problem :)